OpenStudy (anonymous):

PLEASE HELP Will give Medal! ;) Which is the equation of a line that is perpendicular to 8x=6y=-5 A) y=2/3x-3/4 B) y=3/4x-3/4 C) y=4x+6 D) y=6x-4

OpenStudy (perl):

8x + 6y = -5 , is that the equation?

OpenStudy (anonymous):

yes it is

OpenStudy (perl):

ok first solve for y ,

OpenStudy (perl):

after you solve y = mx + b, the slope of the perpindicular line will be -1/m

OpenStudy (anonymous):

i think it y=-8/6-5 but i dont know im terrible

OpenStudy (perl):

almost

OpenStudy (anonymous):

y=-8/6x-5*

OpenStudy (anonymous):

ive been trying to do this on my own but i cant

OpenStudy (perl):

8x + 6y = -5 6y = -8x -5 y = -8/6x - 5/6

OpenStudy (perl):

or in lowest terms y = -4/3x - 5/6

OpenStudy (anonymous):

So how do i find the perpendicular line?

OpenStudy (perl):

after you solve y = mx + b, the slope of the perpindicular line will be -1/m

OpenStudy (perl):

the negative reciprocal

OpenStudy (anonymous):

so the slope would be 3/4x?